# Permissioned LLMs: Enforcing Access Control in Large Language Models

### v1

@@ Abstract
In enterprise settings, organizational data is segregated, siloed and carefully protected by elaborate access control frameworks. These access control structures can completely break down if an LLM fine-tuned on the siloed data serves requests, for downstream tasks, from individuals with disparate access privileges. We propose Permissioned LLMs (PermLLM), a new class of LLMs that superimpose the organizational data access control structures on query responses they generate. We formalize abstractions underpinning the means to determine whether access control enforcement happens correctly over LLM query responses. Our formalism introduces the notion of a relevant response that can be used to prove whether a PermLLM mechanism has been implemented correctly. We also introduce a novel metric, called access advantage, to empirically evaluate the efficacy of a PermLLM mechanism. We introduce three novel PermLLM mechanisms that build on Parameter Efficient Fine-Tuning to achieve the desired access control. We furthermore present two instantiations of access advantage–(i) Domain Distinguishability Index (DDI) based on Membership Inference Attacks, and (ii) Utility Gap Index (UGI) based on LLM utility evaluation. We demonstrate the efficacy of our PermLLM mechanisms through extensive experiments on four public datasets (GPQA, RCV1, SimpleQA, and WMDP), in addition to evaluating the validity of DDI and UGI metrics themselves for quantifying access control in LLMs.
@@ 1 Introduction
This paper focuses on the access control problem for LLMs when they are tuned on data coming from a multitude of data silos. The challenge here is to guarantee that users who do not have access to specific data silos cannot retrieve information from those silos by sending carefully crafted queries to the LLMs tuned on data from those silos. A recent work [12] took an initial step in this direction, but lacks the formal framework to evaluate the access control. Moreover they only explore one type of mechanism. As a security problem, access control is a zero-sum game, hence probabilistic solutions (e.g. Differential Privacy [9]) are not satisfactory.
@@
In this paper, we comprehensively study the problem of access control in LLM fine-tuning. More specifically: (i) We formalize the notion of access control mechanism in LLMs in terms of the relevance of responses generated by an LLM to the raw data the users have access to. We use the notion of security domains in our formalism. Our formalism of response relevance can be used to prove correctness of access control mechanisms. We also propose a novel metric called access advantage that helps us empirically quantify the effectiveness of an access control mechanism on LLMs (§ 2). (ii) We present three new PermLLM fine-tuning mechanisms (see Figure 1), based on Parameter Efficient Fine-Tuning (PEFT) [17, 41] (§ 3). (iii) We introduce two novel instances of our access advantage metric, Domain Distinguishability Index (DDI) and Utility Gap Index (UGI), as tools to audit access control enforcement via an adversarial gaming setting (§ 4). (iv) We empirically evaluate our access control mechanisms on two LLMs (Mistral-0.1-7B and Llama-3.1-8B) using four different data sets: GPQA [32], RCV1 [21], SimpleQA [40], and WMDP [23] (§ 5). Our evaluation shows the effectiveness of our access advantage metrics in assessing whether a proposed access control mechanism for LLMs is enforcing data protection correctly.
@@ 2.2 Definitions
The above definition of relevant response helps us formally determine if a proposed access control mechanism $\mathcal{M}$ is algorithmically correct. We however require an empirically quantifiable metric to determine if the implementation (and the algorithm by extension) of $\mathcal{M}$ is correct. This is particularly important for auditing. To that end, we propose a new metric called response relevance score, $relv(f_{D}^{\mathcal{M}}(q),S_{u})$, which quantifies the information gained on data in the domain set $S_{u}$ by observing responses to queries generated using model parameters $W_{S_{u}}$ affected by domains of $S_{u}$. $relv$ is expected to be higher when $q\sim\mathcal{D}_{S_{u}}$ (i.e., $q$ is related to domain set $S_{u}$), compared to when $q\not\sim\mathcal{D}_{S_{u}}$.
We restrict the domain of $relv$ to the real number interval $[0,1]$, where $1$ is the best expected score for relevance. $relv$ itself can be represented by another empirical metric such as prediction accuracy, or logits for the expected response. However, given that LLMs (and ML models in general) are generalization engines, in practice we expect $relv$ to be less than $1$. This restriction can be effectively addressed by measuring $relv$ for domains that the user has access to and comparing it to $relv$ for domains that the user does not have access to. We call this the access advantage.
@@ 2.3 Auditing Access Control
    $\mathcal{A}$ concludes the access control mechanism is correctly implemented if the access advantage $relv(f_{D}^{\mathcal{M}}(q),S_{u})\circleddash relv(f_{D}^{\mathcal{M}}(q),S_{v})\geq\alpha$.
@@ 4.1 Metric 1: Domain Distinguishability Index (DDI)
For a domain universe $\mathbb{S}$ consisting of $n$ security domains, let $f_{D}^{\mathcal{M}}$ denote the PermLLM trained on data $D$ from all security domains with access control mechanism $\mathcal{M}$. For each ordered pair of domain sets $(S_{i}\subseteq\mathbb{S},S_{j}\subseteq\mathbb{S})$ with no overlap (i.e.,$S_{i}\cap S_{j}=\phi$), let $O^{(S_{i},S_{j})}=O(f_{D}^{\mathcal{M}}(q)|S_{i},f_{D}^{\mathcal{M}}(q)|S_{j});\forall q\sim\mathcal{D}_{S_{i}}$ be the output of a membership inference oracle $O$. For a given membership inference metric $m(\cdot)$, the DDI is defined as: $\mathrm{DDI}(m)=\mathbb{E}_{S_{i}\subseteq\mathbb{S},S_{j}\subseteq\mathbb{S}}\big{[}m\bigl{(}O^{(S_{i},S_{j})}\bigr{)}\big{]}$, where $\mathbb{E}$ is the expectation over all domain sets.
@@
We use AUC-ROC and TPR@$(low)$FPR, as instantiations of DDI, where higher scores indicate stronger enforcement, as $S_{i}$-specific responses become more distinguishable. See Appendices E.1 and E.2 for details on MIA evaluation metrics and an overview of existing MIAs against LLMs.
@@ 4.2 Metric 2: Utility Gap Index (UGI)
Let $U(\cdot)$ be a chosen utility metric and for a domain set pair $(S_{i}\subseteq\mathbb{S},S_{j}\subseteq\mathbb{S})$ with no overlap (i.e.,$S_{i}\cap S_{j}=\phi$), $\mathrm{UtilityGap}^{(S_{i},S_{j})}(U)=|U(f_{D}^{\mathcal{M}}(q)|S_{i})-U(f_{D}^{\mathcal{M}}(q)|S_{j})|;\forall q\sim\mathcal{D}_{S_{i}}$. The UGI aggregates utility gaps across all ordered domain set pairs: $\Delta_{U}\;=\;\mathbb{E}_{S_{i}\subseteq\mathbb{S},S_{j}\subseteq\mathbb{S}}\big{[}\mathrm{UtilityGap}^{(S_{i},S_{j})}(U)\big{]}$, where $\mathbb{E}$ is the expectation over all domain sets.
@@ 5 Experimental Evaluation
For our experiments, we fine-tune Llama-3.1-8B and Mistral-0.1-7B pretrained models on four datasets covering multiple distinct security domains (henceforth called domains), where we fine-tune a separate LoRA adapter for each domain. Details about the model hyperparameters can be found in Appendix § D.1. The data sets we use in our experiments are WMDP [23], GPQA [32], SimpleQA [40], and RCV1 [21]. Table 1 shows the brief data set details. More details on the data sets and generalization gaps can be found in Appendix § D.2. Appendix § D.3 discusses the utility evaluation of all our models.
@@
WMDP GPQA SimpleQA RCV1
Data Set Size (Train / Test) 2936 / 732 360 / 88 4089 / 1018 45622 / 22811
Number of Security Domains 3 3 10 4
@@ 5.1 Evaluating Access Control
Our approach achieves comparable model utility to existing approaches of fine-tuning (see discussion in § D.3), in addition to providing access control. Here we will empirically evaluate the effectiveness of our access control using a suite of metrics. We will first consider the case where the user has access to only one domain and report the access control results in § 5.1.1. Next in § 5.1.2, we will consider the case where the user has access to multiple domains.
@@ 5.1.1 Single Active Domain
Theoretically, $\Delta_{U}$ may reach 1.0, but empirically we observe much smaller—yet substantial—access advantage gaps (Figure 2). These gaps are significantly impacted by domain distributions and the strictness of the scoring metric. For example, SimpleQA exhibits the largest UGIs (up to $\Delta_{blue}\!\approx\!0.50$ and $\Delta_{acc}\!\approx\!0.50$) because it has the highest number of distinct domains (10 in total). Moreover, we observe that $\Delta_{bleu}$ and $\Delta_{acc}$ have the largest values as these metrics look for verbatim pattern matches, thus requiring the model to memorize the nuances in the target domain. On the other hand, $\Delta_{bleurt}$ and $\Delta_{bert}$ look for approximate similarities, and hence are impacted by the similarities across the domains. This suggests that the verbatim matching metrics, $\Delta_{bleu}$ and $\Delta_{acc}$, are better model utility metrics for measuring access advantage compared to the similarity based metrics $\Delta_{bleurt}$ and $\Delta_{bert}$. For large data sets like RCV1, all the metrics achieve similar values as the model begins to generalize more. While these values are not close to 1, they still provide credence to the fact that the domains are different and our access control protocol works as expected due to the utility gaps. The access advantage threshold $\alpha$ is dependent on the type of utility metric: verbatim matching metrics $\Delta_{bleu}$ and $\Delta_{acc}$ have higher threshold than similarity based metrics $\Delta_{bleurt}$ and $\Delta_{bert}$. For $\Delta_{acc}$ metric, $\alpha>0.2$ is sufficient to infer that access control is happening correctly.
Table 2 shows DDI values obtained from a suite of state-of-the-art MIAs. Across domain pairs, the access advantage (distinguishability) scores approach $\alpha~{}=1.0$, indicating that an external auditor can almost perfectly identify the active domain. Hence, even when UGI values fall significantly below $1.0$ because of model generalization, the high DDI values show that access control in Activate still functions as intended. This clearly suggests that DDI is the better method for PermLLM access control efficacy evaluation.
@@
[…]
@@
Mink $0.99\pm 0.01$ $0.94\pm 0.11$ $0.98\pm 0.03$ $1.00\pm 0.00$ $0.98\pm 0.02$ $0.99\pm 0.02$
Mink++ $1.00\pm 0.00$ $1.00\pm 0.01$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
Ref $1.00\pm 0.00$ $0.97\pm 0.04$ $0.99\pm 0.01$ $1.00\pm 0.00$ $0.97\pm 0.05$ $0.99\pm 0.02$
SimpleQA Loss $0.98\pm 0.03$ $0.81\pm 0.34$ $0.90\pm 0.25$ $0.99\pm 0.03$ $0.81\pm 0.32$ $0.92\pm 0.20$
ZLIB $0.98\pm 0.03$ $0.80\pm 0.33$ $0.90\pm 0.23$ $0.99\pm 0.03$ $0.80\pm 0.33$ $0.91\pm 0.20$
Mink $0.99\pm 0.03$ $0.80\pm 0.32$ $0.91\pm 0.21$ $0.99\pm 0.03$ $0.82\pm 0.31$ $0.92\pm 0.20$
Mink++ $0.98\pm 0.03$ $0.81\pm 0.32$ $0.91\pm 0.21$ $0.99\pm 0.03$ $0.82\pm 0.31$ $0.92\pm 0.21$
@@
Mink $1.00\pm 0.01$ $0.94\pm 0.10$ $0.98\pm 0.03$ $0.99\pm 0.01$ $0.88\pm 0.18$ $0.98\pm 0.05$
Mink++ $1.00\pm 0.00$ $0.97\pm 0.05$ $0.99\pm 0.01$ $1.00\pm 0.01$ $0.96\pm 0.06$ $0.99\pm 0.02$
Ref $0.99\pm 0.01$ $0.77\pm 0.28$ $0.99\pm 0.03$ $0.99\pm 0.01$ $0.80\pm 0.28$ $0.98\pm 0.05$
@@ 5.1.2 Multiple Active Domains
As discussed earlier in § 3, we explore three methods of combining knowledge from multiple domains the user has access to: (a) activating all the domain-specific LoRA modules (Activate), (b) merging the LoRA modules (Merge), and (c) training separate LoRA modules on the union of domains and using those for inference (Union). Table 3 reports the UGI ($\Delta_{U}$) for these approaches when the user has access to two domains for all the data sets. We note that WMDP and GPQA have only three security domains, and hence activating any two domains always lead to overlap when calculating $\Delta_{U}$ as per 4.2. For these data sets, we calculate $\Delta_{U}$ on the non-overlapping data. Activate is computationally inexpensive but suffers from considerable utility loss when compared to the previous case of single domain. This is due to the high interference across the multiple domains in the activation space, which is a known issue in the multi-task learning literature [48, 38, 30]. The utility loss suppresses the absolute $\Delta_{U}$ in our experiments. As can be seen in Figure 3, Merge reduces the cross-domain interference, but still suffers from utility loss. Interestingly Merge achieves even lower $\Delta_{U}$ than Activate when combining two domains, as shown in Table 3. Although it quickly outperforms Activate when the user has access to more than two domains, the utility loss due to model merging interference [37, 42, 45, 49] also results in progressive degradation of $\Delta_{U}$ (see Figure 3). Union retains $\Delta_{U}$ even beyond four domains, and hence is the best choice when combining knowledge from several domains. But this comes at the cost of more training-time computation since new domain-specific modules have to be trained for the union of domains, and there could be potential combinatorial blow-up of the number of such combinations. We observe similar results for Mistral-0.1-7B model (see Figure 8 in the appendix).
@@
WMDP $\Delta_{bleurt}$ $0.09\pm 0.01$ $0.07\pm 0.02$ $0.11\pm 0.02$ $0.10\pm 0.02$ $0.08\pm 0.03$ $0.14\pm 0.03$
$\Delta_{bert}$ $0.05\pm 0.01$ $0.03\pm 0.01$ $0.06\pm 0.01$ $0.05\pm 0.01$ $0.04\pm 0.02$ $0.07\pm 0.02$
@@
GPQA $\Delta_{bleu}$ $0.15\pm 0.06$ $0.11\pm 0.06$ $0.51\pm 0.07$ $0.24\pm 0.10$ $0.17\pm 0.10$ $0.62\pm 0.02$
$\Delta_{bleurt}$ $0.10\pm 0.02$ $0.06\pm 0.02$ $0.26\pm 0.03$ $0.14\pm 0.06$ $0.10\pm 0.04$ $0.32\pm 0.02$
$\Delta_{bert}$ $0.07\pm 0.02$ $0.04\pm 0.03$ $0.18\pm 0.02$ $0.11\pm 0.04$ $0.08\pm 0.03$ $0.21\pm 0.02$
$\Delta_{acc}$ $0.09\pm 0.04$ $0.05\pm 0.02$ $0.31\pm 0.08$ $0.16\pm 0.07$ $0.08\pm 0.07$ $0.51\pm 0.04$
@@
RCV1 $\Delta_{bleu}$ $0.05\pm 0.03$ $0.04\pm 0.02$ $0.16\pm 0.09$ $0.04\pm 0.02$ $0.01\pm 0.03$ $0.19\pm 0.10$
$\Delta_{bleurt}$ $0.11\pm 0.01$ $0.07\pm 0.03$ $0.22\pm 0.08$ $0.08\pm 0.01$ $0.03\pm 0.04$ $0.22\pm 0.08$
$\Delta_{bert}$ $0.08\pm 0.01$ $0.06\pm 0.02$ $0.16\pm 0.04$ $0.06\pm 0.01$ $0.03\pm 0.05$ $0.18\pm 0.06$
$\Delta_{acc}$ $0.03\pm 0.01$ $0.04\pm 0.04$ $0.24\pm 0.14$ $0.02\pm 0.02$ $0.01\pm 0.03$ $0.26\pm 0.15$

### v2

@@ Abstract
In enterprise settings, organizational data is segregated, siloed and carefully protected by elaborate access control frameworks. These access control structures can completely break down if an LLM fine-tuned on the siloed data serves requests, for downstream tasks, from individuals with disparate access privileges. We propose Permissioned LLMs (PermLLM), a new class of LLMs that superimpose the organizational data access control structures on query responses they generate. We formalize abstractions underpinning the means to determine whether access control enforcement happens correctly over LLM query responses. Our formalism introduces the notion of a relevant response that can be used to prove whether a PermLLM mechanism has been implemented correctly. We also introduce a novel metric, called access advantage, to empirically evaluate the efficacy of a PermLLM mechanism. We introduce three novel PermLLM mechanisms that build on Parameter Efficient Fine-Tuning to achieve the desired access control. We furthermore present two instantiations of access advantage–(i) Domain Distinguishability Index (DDI) based on Membership Inference Attacks, and (ii) Utility Gap Index (UGI) based on LLM utility evaluation. We demonstrate the efficacy of our PermLLM mechanisms through extensive experiments on five public datasets (GPQA, RCV1, SimpleQA, WMDP, and PubMedQA), in addition to evaluating the validity of DDI and UGI metrics themselves for quantifying access control in LLMs.
@@ 1 Introduction
This paper focuses on the access control problem for LLMs when they are tuned on data coming from a multitude of data silos. The challenge here is to guarantee that users who do not have access to specific data silos cannot retrieve information from those silos by sending carefully crafted queries to the LLMs tuned on data from those silos. A recent work [12] took an initial step in this direction, but lacks the formal framework to evaluate the access control. Moreover they only explore one type of mechanism. As a security problem, access control is absolute– you either achieve access control or not, hence probabilistic solutions (e.g. Differential Privacy [9]) are not satisfactory.
@@
In this paper, we comprehensively study the problem of access control in LLM fine-tuning. More specifically: (i) We formalize the notion of access control mechanism in LLMs in terms of the relevance of responses generated by an LLM to the raw data the users have access to. We use the notion of security domains in our formalism. Our formalism of response relevance can be used to prove correctness of access control mechanisms. We also propose a novel metric called access advantage that helps us empirically quantify the effectiveness of an access control mechanism on LLMs (§ 2). (ii) We present three new PermLLM fine-tuning mechanisms (see Figure 1), based on Parameter Efficient Fine-Tuning (PEFT) [17, 42] (§ 3). (iii) We introduce two novel instances of our access advantage metric, Domain Distinguishability Index (DDI) and Utility Gap Index (UGI), as tools to audit access control enforcement via an adversarial gaming setting (§ 4). (iv) We empirically evaluate our access control mechanisms on two LLMs (Mistral-0.1-7B and Llama-3.1-8B) using five different data sets: GPQA [33], RCV1 [22], SimpleQA [41], WMDP [24], and PubMedQA [20] (§ 5). Our evaluation shows the effectiveness of our access advantage metrics in assessing whether a proposed access control mechanism for LLMs is enforcing data protection correctly.
@@ 2.2 Definitions
The above definition of relevant response helps us formally determine if a proposed access control mechanism $\mathcal{M}$ is algorithmically correct. We however require an empirically quantifiable metric to determine if the implementation (and the algorithm by extension) of $\mathcal{M}$ is correct. This is particularly important for auditing. To that end, we propose a new metric called response relevance score, $relv(f_{D}^{\mathcal{M}}(q),S_{u})$, which quantifies the information gained on data in the domain set $S_{u}$ by observing responses to queries generated using model parameters $W_{S_{u}}$ affected by domains of $S_{u}$. relvrelv is expected to be higher when $q\sim\mathcal{D}_{S_{u}}$ (i.e., $q$ is related to domain set $S_{u}$), compared to when $q\not\sim\mathcal{D}_{S_{u}}$.
We restrict the domain of relvrelv to the real number interval $[0,1]$, where $1$ is the best expected score for relevance. relvrelv itself can be represented by another empirical metric such as prediction accuracy, or logits for the expected response. However, given that LLMs (and ML models in general) are generalization engines, in practice we expect relvrelv to be less than $1$. This restriction can be effectively addressed by measuring relvrelv for domains that the user has access to and comparing it to relvrelv for domains that the user does not have access to. We call this the access advantage.
@@ 2.3 Auditing Access Control
    $\mathcal{A}$ concludes the access control mechanism is correctly implemented if the access advantage $|relv(f_{D}^{\mathcal{M}}(q),S_{u})\circleddash relv(f_{D}^{\mathcal{M}}(q),S_{v})|\geq\alpha$.
@@ 4.1 Metric 1: Domain Distinguishability Index (DDI)
[…]
We use AUC-ROC and TPR@(low)(low)FPR, as instantiations of DDI, where higher scores indicate stronger enforcement, as $S_{i}$-specific responses become more distinguishable. See Appendices F.1 and F.2 for details on MIA evaluation metrics and an overview of existing MIAs against LLMs.
@@ 4.2 Metric 2: Utility Gap Index (UGI)
[…]
@@ 5 Experimental Evaluation
For our experiments, we fine-tune Llama-3.1-8B and Mistral-0.1-7B pretrained models on five datasets covering multiple distinct security domains (henceforth called domains), where we fine-tune a separate LoRA adapter for each domain. Details about the model hyperparameters can be found in Appendix § D.1. The data sets we use in our experiments are WMDP [24], GPQA [33], SimpleQA [41], RCV1 [22], and PubMedQA [20]. Table 1 shows the brief data set details. More details on the data sets and generalization gaps can be found in Appendix § D.2. Appendix § D.3 discusses the utility evaluation of all our models.
@@
WMDP GPQA SimpleQA RCV1 PubMedQA
Data Set Size (Train / Test) 2936 / 732 360 / 88 4089 / 1018 45622 / 22811 200000 / 11269
Number of Security Domains 3 3 10 4 10
@@ 5.1 Evaluating Access Control
Our approach achieves comparable model utility to existing approaches of fine-tuning (see discussion in § D.3), in addition to providing access control. Here we will empirically evaluate the effectiveness of our access control using a suite of metrics. We will first consider the case where the user has access to only one domain and report the access control results in § 5.1.1. Next in § 5.1.2, we will consider the case where the user has access to multiple domains. For comparison, we also include an evaluation of a prompt-based access control baseline in Appendix E but find it to be ineffective.
@@ 5.1.1 Single Active Domain
Theoretically, $\Delta_{U}$ may reach 1.0, but empirically we observe much smaller—yet substantial—access advantage gaps for four of the data sets (Figure 2). These gaps are significantly impacted by domain distributions and the strictness of the scoring metric. For example, SimpleQA exhibits the largest UGIs (up to $\Delta_{blue}\!\approx\!0.50$ and $\Delta_{acc}\!\approx\!0.50$) because it has the highest number of distinct domains (10 in total). Moreover, we observe that $\Delta_{bleu}$ and $\Delta_{acc}$ have the largest values as these metrics look for verbatim pattern matches, thus requiring the model to memorize the nuances in the target domain. On the other hand, $\Delta_{bleurt}$ and $\Delta_{bert}$ look for approximate similarities, and hence are impacted by the similarities across the domains. This suggests that the verbatim matching metrics, $\Delta_{bleu}$ and $\Delta_{acc}$, are better model utility metrics for measuring access advantage compared to the similarity based metrics $\Delta_{bleurt}$ and $\Delta_{bert}$. For large data sets like RCV1, all the metrics achieve similar values as the model begins to generalize more. While these values are not close to 1, they still provide credence to the fact that the domains are different and our access control protocol works as expected due to the utility gaps. The access advantage threshold $\alpha$ is dependent on the type of utility metric: verbatim matching metrics $\Delta_{bleu}$ and $\Delta_{acc}$ have higher threshold than similarity based metrics $\Delta_{bleurt}$ and $\Delta_{bert}$. For $\Delta_{acc}$ metric, $\alpha>0.2$ is sufficient to infer that access control is happening correctly. PubMedQA is an exception where $\Delta_{U}$ values are close to zero; this is because the security domains are artificially obtained via k-means and hence have the same underlying data distribution.
Table 2 shows DDI values obtained from a suite of state-of-the-art MIAs. Across domain pairs, the access advantage (distinguishability) scores approach $\alpha~=1.0$, indicating that an external auditor can almost perfectly identify the active domain (even when the domain distributions are similar as in the case of PubMedQA). Hence, even when UGI values fall significantly below $1.0$ because of model generalization, the high DDI values show that access control in Activate still functions as intended. This clearly suggests that DDI is the better method for PermLLM access control efficacy evaluation.
@@
Table 2: DDI values with $m\in\{\text{AUC–ROC},\text{TPR@1\%FPR},\text{TPR@5\%FPR}\}$ for the different MIAs. Mink and Mink++ are run with hyperparameter $k=10\%$. Entries are reported as $mean\pm std$ across security domains.
@@
Mink $0.99\pm 0.01$ $0.94\pm 0.11$ $0.98\pm 0.03$ $1.00\pm 0.00$ $8\pm 0.02 0.99±0$$9\pm 0.02$
Mink++ $\pm 0.00 1.00±0.01$$\pm 0.01$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
Ref $1.00\pm 0.00$ $7\pm 0.04 0.99±0$$9\pm 0.01$ $1.00\pm 0.00$ $7\pm 0.05 0.99±0$$9\pm 0.02$
SimpleQA Loss $0.98\pm 0.03$ $0.81\pm 0.34$ $0\pm 0.25 0.99±0$$9\pm 0.03$ $0.81\pm 0.32$ $0.92\pm 0.20$
ZLIB $0.98\pm 0.03$ $0.80\pm 0.33$ $0\pm 0.23 0.99±0$$9\pm 0.03$ $0.80\pm 0.33$ $0.91\pm 0.20$
Mink $0.99\pm 0.03$ $0.80\pm 0.32$ $1\pm 0.21 0.99±0$$9\pm 0.03$ $0.82\pm 0.31$ $0.92\pm 0.20$
Mink++ $0.98\pm 0.03$ $0.81\pm 0.32$ $1\pm 0.21 0.99±0$$9\pm 0.03$ $0.82\pm 0.31$ $0.92\pm 0.21$
@@
Mink $1.00\pm 0.01$ $0.94\pm 0.10$ $8\pm 0.03 0.99±0$$9\pm 0.01$ $0.88\pm 0.18$ $0.98\pm 0.05$
Mink++ $1.00\pm 0.00$ $7\pm 0.05 0.99±0$$9\pm 0.01$ $1.00\pm 0.01$ $6\pm 0.06 0.99±0$$9\pm 0.02$
Ref $0.99\pm 0.01$ $0.77\pm 0.28$ $9\pm 0.03 0.99±0$$9\pm 0.01$ $0.80\pm 0.28$ $0.98\pm 0.05$
PubMedQA Loss $0.81\pm 0.07$ $0.16\pm 0.11$ $0.36\pm 0.15$ $0.95\pm 0.03$ $0.51\pm 0.21$ $0.75\pm 0.14$
ZLIB $0.77\pm 0.07$ $0.10\pm 0.05$ $0.30\pm 0.13$ $0.88\pm 0.05$ $0.32\pm 0.17$ $0.57\pm 0.15$
Mink $0.86\pm 0.05$ $0.25\pm 0.12$ $0.48\pm 0.15$ $0.98\pm 0.01$ $0.75\pm 0.14$ $0.91\pm 0.07$
Mink++ $0.90\pm 0.02$ $0.31\pm 0.08$ $0.57\pm 0.08$ $0.99\pm 0.01$ $0.93\pm 0.07$ $0.98\pm 0.02$
Ref $1.00\pm 0.00$ $0.98\pm 0.02$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
@@ 5.1.2 Multiple Active Domains
As discussed earlier in § 3, we explore three methods of combining knowledge from multiple domains the user has access to: (a) activating all the domain-specific LoRA modules (Activate), (b) merging the LoRA modules (Merge), and (c) training separate LoRA modules on the union of domains and using those for inference (Union). Table 3 reports the UGI ($\Delta_{U}$) for these approaches when the user has access to two domains for all the data sets. We note that WMDP and GPQA have only three security domains, and hence activating any two domains always lead to overlap when calculating $\Delta_{U}$ as per 4.2. For these data sets, we calculate $\Delta_{U}$ on the non-overlapping data. Activate is computationally inexpensive but suffers from considerable utility loss when compared to the previous case of single domain. This is due to the high interference across the multiple domains in the activation space, which is a known issue in the multi-task learning literature [49, 39, 31]. The utility loss suppresses the absolute $\Delta_{U}$ in our experiments. As can be seen in Figure 3, Merge reduces the cross-domain interference, but still suffers from utility loss. Interestingly Merge achieves even lower $\Delta_{U}$ than Activate when combining two domains, as shown in Table 3. Although it quickly outperforms Activate when the user has access to more than two domains, the utility loss due to model merging interference [38, 43, 46, 50] also results in progressive degradation of $\Delta_{U}$ (see Figure 3). Union retains $\Delta_{U}$ even beyond four domains, and hence is the best choice when combining knowledge from several domains. But this comes at the cost of more training-time computation since new domain-specific modules have to be trained for the union of domains, and there could be potential combinatorial blow-up of the number of such combinations. As with the single active domain case, we observe close to zero utility gap on PubMedQA as the domains share the same data distribution. We observe similar results for Mistral-0.1-7B model (see Figure 9 in the appendix).
@@
WMDP $\Delta_{bleurt}$ $09\pm 0.01 0.0$$07\pm 0.02$ $0.11\pm 0.02$ $10\pm 0.02 0.0$$08\pm 0.03$ $0.14\pm 0.03$
$\Delta_{bert}$ $05\pm 0.01 0.0$$3\pm 0.01 0.0$$6\pm 0.01 0.0$$5\pm 0.01 0.0$$4\pm 0.02 0.0$$07\pm 0.02$
@@
GPQA $\Delta_{bleu}$ $5\pm 0.06 0.11±0$$1\pm 0.06$ $0.51\pm 0.07$ $0.24\pm 0.10$ $0.17\pm 0.10$ $0.62\pm 0.02$
$\Delta_{bleurt}$ $10\pm 0.02 0.0$$06\pm 0.02$ $0.26\pm 0.03$ $0.14\pm 0.06$ $0.10\pm 0.04$ $0.32\pm 0.02$
$\Delta_{bert}$ $07\pm 0.02 0.0$$04\pm 0.03$ $8\pm 0.02 0.11±0$$1\pm 0.04$ $0.08\pm 0.03$ $0.21\pm 0.02$
$\Delta_{acc}$ $09\pm 0.04 0.0$$05\pm 0.02$ $0.31\pm 0.08$ $16\pm 0.07 0.0$$08\pm 0.07$ $0.51\pm 0.04$
@@
RCV1 $\Delta_{bleu}$ $05\pm 0.03 0.0$$04\pm 0.02$ $16\pm 0.09 0.0$$4\pm 0.02 0.0$$01\pm 0.03$ $0.19\pm 0.10$
$\Delta_{bleurt}$ $11\pm 0.01 0.0$$07\pm 0.03$ $22\pm 0.08 0.0$$8\pm 0.01 0.0$$03\pm 0.04$ $0.22\pm 0.08$
$\Delta_{bert}$ $08\pm 0.01 0.0$$06\pm 0.02$ $16\pm 0.04 0.0$$6\pm 0.01 0.0$$03\pm 0.05$ $0.18\pm 0.06$
$\Delta_{acc}$ $03\pm 0.01 0.0$$04\pm 0.04$ $24\pm 0.14 0.0$$2\pm 0.02 0.0$$01\pm 0.03$ $0.26\pm 0.15$
PubMedQA $\Delta_{bleu}$ $1\pm 0.00 0.00±0$$\pm 0.00 0.0$$1\pm 0.00 0.0$$1\pm 0.00 0.0$$0\pm 0.00 0.0$$01\pm 0.01$
$\Delta_{bleurt}$ $1\pm 0.00 0.00±0$$\pm 0.00 0.0$$1\pm 0.00 0.0$$1\pm 0.00 0.0$$0\pm 0.00 0.0$$01\pm 0.01$
$\Delta_{bert}$ $1\pm 0.00 0.00±0$$\pm 0.00 0.0$$1\pm 0.00 0.0$$1\pm 0.00 0.0$$0\pm 0.00 0.0$$01\pm 0.00$
$\Delta_{acc}$ - - - - - -